A sous chef in a fancy restaurant boasts of cooking steak perfectly 96% of the time. However one evening, he has already undercooked 3 out of the 20 orders and the orders have been returned to the kitchen. How likely is this? Assuming an a of 0.01, is there sufficient evidence from this evening to suggest that his claim is true?
Answer #1

Let X denote the number of steak that is not properly cooked. Then X ~ Bin(n-20. p= 0.04)

So, PlX-z) _ ()0.0(1-0.04)»,-0,1, 2, 3, 0.041 0.04), r 0,1, 2,3, 20

Required probability = PlX-3)-(#90043(1-004 20-3 0.0364

Let p and P denote the sample and population proportion of steak that is not properly cooked respectively. Then,

H0: P = 0.04

Ha: Pメ0.04

Now, under H0, P(1- P) 0.041 0.04) N(0.04

Test statistic value = p- P 0.04 0.04(1-0.04) 7t

Since this is two-tailed test, so p-value is given by: 2 × P(Z > 2.51)-2 × |1 _ P(Z < 2.51)| 2 x |1-0.9940-0.012

Since p-value = 0.012 is greater than 0.01, so we do not have sufficient evidence to reject H0. Thus we say, P = 0.04 i.e we say his claim is true.